6.0 grams hydrocarbon sample were made to react with oxygen gas to produce 17.6 grams of carbon dioxide according to the equation:

CxHy(g) + O2(g) ---------- CO2(g) + H20(l) where x = subscript for C, y = subscript for H

a. How many grams of C are there in 17.6 grams of CO2?

b. How many grams of C are there in the hydrocarbon sample?

c. How many moles of C are there in the hydrocarbon sample?

d. How many grams of H are there in the hydrocarbon sample?

e. How many moles of H are there in the hydrocarbon sample?

f. What is the empirical formula of the hydrocarbon sample.

g. If the Molar Mass of the hydrocarbon sample is 60 g/mole, what is the molecular formula of the hydrocarbon sample?

h. What is the % composition of the hydrocarbon sample (%C, %H)
Answer #1

a. How many grams of C are there in 17.6 grams of CO2?

there are 12 g of C per 44 g of CO2 ( molar mass ratios)

therefore

mass of C = 12/44*17.6 = 4.8 g are Carbon

b. How many grams of C are there in the hydrocarbon sample?

Clearly, all Carbon comes form the hydrocarbon

so

mass of C in HC =  mass of Carbon in CO2

mass of C in HC = 4.8 g of Carbon in Hydrocarbon

c. How many moles of C are there in the hydrocarbon sample?

mol = mass/MW

mol = 4.8/12 = 0.4 mol of C

d. How many grams of H are there in the hydrocarbon sample?

mass of H in HC = total mass - mass of carbon = 6 - 4.8 = 1.2 g of Hydrogen